a secretary types 5 letters and 5 envelopes buts the letter into the envelop at random.what is the probability that at least 1 letter is put into its correct envelop
Answer by stanbon(75887)

a secretary types 5 letters and 5 envelopes puts the letters into the envelops at random.what is the probability that at least 1 letter is put into its correct envelope
------
P(at least one correct) = 1 - p(all incorrect)
----
P(all incorrect) = 1/2^5 = 1/32
---
P(at least one correct) = 1 - 1/32 = 31/32
====================
Cheers,
Stan H.
============